a motorcyclist is moving at ( N+30) m/s when the rider applies the brakes giving the motorcyclist a constant deceleration of 5 meter per square . what distance does the motorcycle travel from the instant braking begins until the motorcyclist stops ?


1
Expert's answer
2021-01-24T16:41:27-0500
"v^2=2as\\30^2=2(5)s\\s=90\ m"